# REASSEMBLE: A Multimodal Dataset for Contact-rich Robotic Assembly and Disassembly

### v1

@@ Abstract
Robotic manipulation remains a core challenge in robotics, particularly for contact-rich tasks such as industrial assembly and disassembly. Existing datasets have significantly advanced learning in manipulation but are primarily focused on simpler tasks like object rearrangement, falling short of capturing the complexity and physical dynamics involved in assembly and disassembly. To bridge this gap, we present REASSEMBLE (Robotic assEmbly disASSEMBLy datasEt), a new dataset designed specifically for contact-rich manipulation tasks. Built around the NIST Assembly Task Board 1 benchmark, REASSEMBLE includes four actions (pick, insert, remove, and place) involving 17 objects. The dataset contains 4,551 demonstrations, of which 4,035 were successful, spanning a total of 781 minutes. Our dataset features multi-modal sensor data including event cameras, force-torque sensors, microphones, and multi-view RGB cameras. This diverse dataset supports research in areas such as learning contact-rich manipulation, task condition identification, action segmentation, and more. We believe REASSEMBLE will be a valuable resource for advancing robotic manipulation in complex, real-world scenarios. The dataset is publicly available on our project website: https://dsliwowski1.github.io/REASSEMBLE_page/.
@@ I Introduction
Robot learning and deep learning, in general, has seen remarkable progress in recent years, driven by the increasing availability of high-quality, large-scale datasets used to train models [2]. These datasets have enabled state-of-the-art algorithms to achieve impressive performance across a wide range of tasks, including motion planning [3, 4, 5], task planning [6, 7, 8], locomotion [9], and anomaly detection [6, 10]. Previous research on robot learning mainly focused on simpler manipulation tasks, such as object rearrangement, picking, and placing. These tasks typically involve short-horizon objectives and limited interaction with the environment. In contrast, more complex challenges, such as long-horizon and contact-rich manipulation tasks, remain less explored. Examples include object assembly and disassembly, which require understanding of interaction dynamics and the ability to plan and execute precise, goal-oriented action sequences. Recent studies, including [11], have highlighted the limitations of current state-of-the-art algorithms to address such tasks. A significant contributing factor to these limitations is the lack of datasets explicitly designed for long-horizon and contact-rich manipulation scenarios. Without targeted data, models struggle to generalize and perform effectively in these challenging environments. Although industries often rely on specialized tools and structured environments to address complex manipulation tasks, robots operating in homes or unstructured environments face significantly greater challenges. These settings demand versatility and adaptability, as tasks often involve a wide variety of objects, irregular conditions, and unpredictable scenarios. Consider a household robot tasked with assembling furniture, such as threading screws, inserting pegs, or aligning interlocking parts. Unlike industrial setups with precise tooling, these tasks require the robot to handle uncertainties in alignment, material tolerances, and force application. Such challenges are even more pronounced in unstructured environments, where small deviations can lead to task failure.
@@
Although much of the focus has traditionally been on enabling robots to assemble products, there exists an equally critical need to develop robots capable of disassembling objects. Disassembly is a cornerstone of sustainability, particularly in advancing the circular economy. According to the United Nations Sustainable Development Goal 12, global demand for natural resources is projected to triple by 2050—far exceeding what the planet can sustainably provide under current consumption patterns. This underscores the urgent need for innovative solutions to improve resource efficiency and increase recycling rates. Robots capable of disassembling objects can play a transformative role in this effort. For instance, they could be deployed to efficiently disassemble complex objects such as cars, electronic devices, or EV batteries, which are currently dismantled manually in labor-intensive and hazardous processes. Despite the critical importance of disassembly in recycling, automation in this sector remains limited. Developing robust, adaptive agents capable of safely and efficiently disassembling products could drastically reduce waste, conserve resources, and mitigate the environmental impact of human consumption.
@@
To bridge the gap between these pressing challenges, we introduce REASSEMBLE, a comprehensive dataset tailored to long-horizon and contact-rich manipulation tasks. While it draws inspiration from industrial assembly and disassembly, REASSEMBLE’s impact extends beyond industrial settings, offering opportunities to advance robotic capabilities for unstructured and everyday environments. Built upon the standardized NIST Assembly Task Boards [1], specifically Task Board #1, REASSEMBLE provides benchmarks for tasks such as gear meshing, peg insertions, electrical connector assemblies, and nut threading, as shown in Figure LABEL:fig:teaser. These tasks mimic the intricate manipulation challenges that robots face in real-world applications. What sets REASSEMBLE apart from other robot manipulation datasets is its focus on multimodal data for holistic learning frameworks. We provide a comparison of commonly used robot learning datasets and their properties in Table I. The REASSEMBLE dataset includes multi-view RGB images and robot proprioceptive data, such as joint states, gripper states, and end-effector poses. Additionally, it incorporates less commonly utilized sensory modalities, such as audio recordings from external and wrist-mounted microphones, as well as six-axis force-torque measurements. Importantly, to the best of our knowledge, REASSEMBLE is the first robot manipulation dataset to include event camera data. Unlike standard RGB cameras, event cameras capture changes in light intensity at each pixel rather than recording entire frames. This low-latency, high-precision information provides critical insights into object movement, enabling models to focus attention effectively and improving robustness in dynamic environments. Furthermore, REASSEMBLE distinguishes itself by providing multi-task annotations, enabling applications across diverse fields such as temporal action segmentation, motion policy learning, and anomaly detection. By offering a rich, multi-modal dataset, REASSEMBLE fosters the development of adaptive and versatile robotic systems capable of tackling the challenges of long-horizon, contact-rich manipulation.
To summarize, our contributions are as follows:
@@
    A multi-modal dataset designed for long-horizon, high-precision, contact-rich manipulation tasks.
@@
    A dataset with multi-task labels to support algorithm development in various robot learning fields, like temporal action segmentation, motion policy learning, and anomaly detection.
@@
REASSEMBLE 4k 4 3 RGB cameras, Robot Proprioception, 1 DAVIS Event camera, 3 Microphones, F/T Sensor Franka Emika Panda Reaserach 3 Haptic teleoperation TAS, MPL, AD, TIL
TABLE I: Datasets comparison. We compare several commonly used datasets based on the number of demonstrations, the number of verbs they contain, the sensors used during data collection, the robotic platform, the data collection method, and the tasks that can be learned from the dataset. The tasks include TAS (Temporal Action Segmentation), MPL (Motion Policy Learning), AD (Anomaly Detection), and TIL (Task Inversion Learning). We use ”n/r” to denote information that is not reported by the works and ”N/A” for cases where the category is not applicable.
@@ II Related Works
Existing datasets, such as the Furniture Benchmark [14], have made progress in addressing long-horizon tasks like furniture assembly. However, they lack the high-quality force-torque data required for the tight tolerances demanded in these applications. Similarly, while datasets like RH20T include manipulation data with audio and force-torque sensor information, they do not address the long-horizon task sequences needed for complex assemblies [16]. Vision-based tactile datasets, such as SPARSH, provide valuable insights into contact-rich manipulation, but they too fall short in capturing the extended temporal dynamics and multimodal sensing required for precise assembly and disassembly tasks [24]. To address these limitations, REASSEMBLE introduces a novel dataset incorporating high-resolution force-torque sensing specifically tailored for low-tolerance, high-precision, and long-horizon tasks.
@@
The increasing prevalence of automation in robotic manipulation tasks highlights the necessity of effective skill assessment, task monitoring, and summarization to enhance system performance and reliability. Temporal action segmentation (TAS) plays a critical role in achieving these objectives by identifying and classifying distinct actions within continuous temporal sequences. Numerous datasets have been developed to support temporal action segmentation [18, 25, 26]. For example, the 50Salads dataset is a widely used benchmark where humans perform salad preparation tasks [18]. This dataset provides temporally labeled actions for long-duration videos, facilitating the training and evaluation of models for action segmentation. These models can learn to segment complex temporal sequences into meaningful subtasks. However, such datasets primarily focus on human activity and often lack relevance to robotic manipulation tasks. Robotics-specific datasets, such as RT-1, Bridge V2, and RoboSet, offer data tailored to specific robotic tasks [27, 13]. While these datasets are valuable for advancing task-specific learning, they are not designed for comprehensive temporal action segmentation across diverse manipulation contexts. Moreover, conventional action segmentation datasets like 50Salads predominantly rely on visual and temporal information, often overlooking critical multimodal data, such as force-torque measurements, which are essential for understanding robotic actions. Therefore, REASSEMBLE also addresses robotic action segmentation and incorporates multimodal data, including visual, force-torque, and temporal streams.
@@
Advancements in large vision-language models (LVLMs), such as PaLM-E, LLaVa, have shown impressive capabilities in reasoning about task conditions by leveraging visual information during robotic task execution [8, 28]. These models provide robust insights into system states, enabling robots to monitor and adapt to dynamic conditions. However, their significant computational requirements make them impractical for real-time applications, where lightweight and efficient models are essential for rapid decision-making. To address these challenges, the ConditionNet was introduced as a focused resource for condition monitoring in robotic systems [21]. While ConditionNet dataset ((Im)PerfectPour) is not aimed at contact-rich manipulation and long-horizon task. Building on these limitations, REASSEMBLE is designed to address the gaps in existing resources. It includes conditioned labeled multi-modal data such as visual, force-torque, and temporal streams, enabling more comprehensive condition monitoring. Crucially, it incorporates failure data to train models that can effectively learn to detect, understand, and respond to failures in real time.
Event cameras are emerging as critical sensors in robotics, providing a fundamentally different data acquisition paradigm compared to traditional RGB cameras [29]. Unlike frame-based cameras that capture static images at fixed intervals, event cameras asynchronously record changes in pixel intensity, resulting in high temporal resolution and low latency. These characteristics make them particularly suited for dynamic environments and fast-changing tasks such as SLAM and locomotion in unstructured settings [30]. Despite the demonstrated utility of event cameras in applications like SLAM and dynamic navigation, to the best of our knowledge, no existing dataset incorporates event camera data for robotic manipulation tasks. This omission represents a significant gap in current research, as event cameras can provide invaluable information for robotic manipulation by capturing motion events with exceptional precision. For instance, during manipulation tasks, event cameras can highlight fine-grained movements, such as the subtle shifts of a robotic gripper or the motion of objects in the environment, even under challenging lighting conditions. One potential application of event camera data in robotic manipulation is the creation of attention masks derived from motion events. These attention masks focus on areas of movement, filtering out irrelevant static regions of the scene. This capability is particularly beneficial for monitoring tasks and identifying critical conditions, such as failures. REASSEMBLE addresses this gap by incorporating event camera data into its multi-modal framework. It provides a unique opportunity to explore the role of event cameras in robotic manipulation, particularly for complex tasks like assembly and disassembly with a NIST Board #1.
@@ III-A REASSEMBLE Tasks
The REASSEMBLE dataset encompasses both assembly and disassembly processes, with the associated actions defined as follows: for assembly, the actions are picking and inserting, whereas for disassembly, the actions are removing and placing. These four actions represent the minimal set required to comprehensively describe the assembly and disassembly tasks. The manual labeling of demonstrations by human operators necessitated limiting the action set to these four categories, thereby reducing cognitive load during annotation. The assembly board contains a total of 20 objects; however, due to constraints in the robot and teleoperation setup, the three smallest bolts were excluded from the dataset (M12, M8, and M4). As a result, the dataset comprises 17 objects. A detailed breakdown of the objects included in the dataset is provided in Figure 3.
@@ III-B Sensors
In addition to the aforementioned sensors, the dataset includes an event camera, a unique feature that sets it apart from similar datasets and provides an advantage in capturing high-speed motion with low latency. The sensor is mounted externally on a tripod in front of the robot. Due to its limited resolution and field-of-view, it is positioned at a distance that allows to capture the majority of the workspace of the robot while still being able to perceive the smaller objects, such as small gears or pegs.
@@ III-C Teleoperation Setup
where $\mathbf{x}_{r}^{\text{init}}\in\mathbb{R}^{3}$ is the initial position of the robot’s end effector. The relative orientation change between the master’s orientation $\mathbf{q}_{m}=[w_{m},\mathbf{v}_{m}^{T}]\in\mathbb{S}^{4}$ and its initial orientation $\mathbf{q}_{m}^{\text{init}}\in\mathbb{S}^{4}$ is calculated as
@@
Given this quaternion difference $\Delta\mathbf{q}_{m}$ and the robot’s initial orientation $\mathbf{q}_{s}^{\text{init}}\in\mathbb{S}^{4}$, the target orientation of the robot end effector is computed as:
@@
The control input $\tau_{\rm{imp}}\in\mathbb{R}^{6}$ for the robot’ Cartesian impedance controller is expressed as
$\tau_{\rm{imp}}=\mathbf{K}_{x}(\boldsymbol{\xi}_{\mathrm{r,cmd}}-\boldsymbol{\xi}_{\mathrm{r}})+\mathbf{D}(\dot{\boldsymbol{\xi}}_{\mathrm{r,cmd}}-\dot{\boldsymbol{\xi}}_{\mathrm{r}}),$ (5)
where $\boldsymbol{\xi}_{\mathrm{r}}\in\mathbb{R}^{6}$ and $\boldsymbol{\xi}_{\mathrm{r,cmd}}\in\mathbb{R}^{6}$ are the current and commanded Cartesian pose of the robot. The matrices $\mathbf{K}_{x}$ and $\mathbf{D}$ represent the stiffness and damping coefficients, respectively, while $\dot{\boldsymbol{\xi}}_{\mathrm{r}}\in\mathbb{R}^{6}$ and $\dot{\boldsymbol{\xi}}_{\mathrm{r,cmd}}\in\mathbb{R}^{6}$ denote the current and commanded task-space velocities.
@@
$V(\boldsymbol{\theta})=\alpha\det\left(\mathbf{J}(\boldsymbol{\theta})\mathbf{J}^{\top}(\boldsymbol{\theta})\right)+\beta\mathrm{JL}(\boldsymbol{\theta})+\gamma\dot{\boldsymbol{\theta}},$ (9)
where $\alpha,\beta,\gamma\in\mathbb{R}$ are weighting coefficients. The determinant term $\det\left(\mathbf{J}(\boldsymbol{\theta})\mathbf{J}^{\top}(\boldsymbol{\theta})\right)$ promotes the avoidance of singularities by maximizing the manipulability measure. The joint limit avoidance term $\mathrm{JL}(\boldsymbol{\theta})$ is expressed as
$\mathrm{JL}(\boldsymbol{\theta})=\tanh\left(a(\boldsymbol{\theta}-\mathbf{c})+b\right)+\tanh\left(-a(\boldsymbol{\theta}-\mathbf{c})+b\right)+2,$ (10)

### v2

@@ Abstract
Robotic manipulation remains a core challenge in robotics, particularly for contact-rich tasks such as industrial assembly and disassembly. Existing datasets have significantly advanced learning in manipulation but are primarily focused on simpler tasks like object rearrangement, falling short of capturing the complexity and physical dynamics involved in assembly and disassembly. To bridge this gap, we present REASSEMBLE (Robotic assEmbly disASSEMBLy datasEt), a new dataset designed specifically for contact-rich manipulation tasks. Built around the NIST Assembly Task Board 1 benchmark, REASSEMBLE includes four actions (pick, insert, remove, and place) involving 17 objects. The dataset contains 4,551 demonstrations, of which 4,035 were successful, spanning a total of 781 minutes. Our dataset features multi-modal sensor data, including event cameras, force-torque sensors, microphones, and multi-view RGB cameras. This diverse dataset supports research in areas such as learning contact-rich manipulation, task condition identification, action segmentation, and task inversion learning. The REASSEMBLE will be a valuable resource for advancing robotic manipulation in complex, real-world scenarios. The dataset is publicly available on our project website [^1] https://tuwien-asl.github.io/REASSEMBLE_page/.
@@ I Introduction
Robot learning, and deep learning more broadly, has made remarkable progress in recent years, driven by the increasing availability of high-quality, large-scale datasets [2]. These datasets have enabled state-of-the-art algorithms to achieve impressive performance across diverse tasks, including motion planning [3, 4, 5], task planning [6, 7, 8], locomotion [9], and anomaly detection [6, 10].
Most previous work in robot learning has focused on relatively simple manipulation tasks, such as object rearrangement, picking, and placing, which typically involve short horizons and limited interaction. In contrast, long-horizon, contact-rich manipulation tasks, such as object assembly and disassembly, remain underexplored. These tasks demand a deeper understanding of interaction dynamics and the ability to plan and execute precise, goal-oriented sequences. Recent work [11] has shown that current algorithms struggle with such tasks, largely due to the lack of datasets tailored for long-horizon, contact-rich scenarios. Without suitable data, models fail to generalize effectively in these challenging settings.
While industrial environments mitigate these challenges through specialized tooling and structure, robots in homes or unstructured settings face greater complexity. Tasks often involve diverse objects, irregular conditions, and unexpected events. For example, a household robot assembling furniture must handle uncertainties in alignment, material tolerances, and force application, without the aid of precise jigs or fixtures. In such cases, small deviations can easily lead to task failure. Many approaches tackle these challenges by decomposing complex actions into simpler skills that can be sequenced to perform long-horizon tasks [12, 13]. For instance, fastening a screw can be broken down into approaching the socket, aligning the screw, and twisting it until secured. This hierarchical learning improves generalization, enabling the reuse of previously learned skills to solve new, unseen tasks [13].
@@
Although much of the focus has traditionally been on enabling robots to assemble products, there exists an equally critical need to develop robots capable of disassembling objects. Disassembly is a cornerstone of sustainability, particularly in advancing the circular economy. According to the United Nations Sustainable Development Goal 12, global demand for natural resources is projected to triple by 2050, far exceeding what the planet can sustainably provide under current consumption patterns. This underscores the urgent need for innovative solutions to improve resource efficiency and increase recycling rates. Robots capable of disassembling objects can play a transformative role in this effort. For instance, they could be deployed to efficiently disassemble complex objects such as cars, electronic devices, or EV batteries, which are currently dismantled manually in labor-intensive and hazardous processes. Despite the critical importance of disassembly in recycling, automation in this sector remains limited. Additionally, this disassembly can be facilitated through task inversion learning (TIL) using assembly demonstrations. Developing robust, adaptive agents capable of safely and efficiently disassembling products could drastically reduce waste, conserve resources, and mitigate the environmental impact of human consumption.
@@
To bridge the gap between these pressing challenges, we introduce REASSEMBLE, a comprehensive dataset tailored to long-horizon and contact-rich manipulation tasks. While it draws inspiration from industrial assembly and disassembly, REASSEMBLE’s impact extends beyond industrial settings, offering opportunities to advance robotic capabilities for unstructured and everyday environments. Built upon the standardized NIST Assembly Task Boards [1], specifically Task Board #1, REASSEMBLE provides benchmarks for tasks such as gear meshing, peg insertions, electrical connector assemblies, and fastening nuts, as shown in Figure LABEL:fig:teaser. These tasks mimic the intricate manipulation challenges that robots face in real-world applications. What sets REASSEMBLE apart from other robot manipulation datasets is its focus on multimodal data for holistic learning frameworks. We provide a comparison of commonly used robot learning datasets and their properties in Table I. The REASSEMBLE dataset includes multi-view RGB images and robot proprioceptive data, such as joint states, gripper states, and end-effector poses. Additionally, it incorporates less commonly utilized sensory modalities, such as audio recordings from external and wrist-mounted microphones, as well as six-axis force-torque measurements. Importantly, REASSEMBLE also includes event camera data. Unlike standard RGB cameras, event cameras capture changes in light intensity at each pixel rather than recording entire frames. This low-latency, high-precision information provides critical insights into object movement, enabling models to focus attention effectively and improving robustness in dynamic environments. Furthermore, REASSEMBLE distinguishes itself by providing multi-task annotations, enabling applications across diverse fields such as hierarchical temporal action segmentation, motion policy learning, and anomaly detection. By offering a rich, multi-modal dataset, REASSEMBLE fosters the development of adaptive and versatile robotic systems capable of tackling the challenges of long-horizon, contact-rich manipulation. To summarize, our contributions are as follows:
@@
    A multi-modal dataset designed for long-horizon, high-precision, contact-rich assembly and disassembly tasks.
@@
    A dataset with multi-task labels to support algorithm development in various robot learning fields, like hierarchical temporal action segmentation, motion policy learning, and anomaly detection.
@@
REASSEMBLE 4k 4 actions and 9 skills 3 RGB cameras, Robot Proprioception, 1 DAVIS Event camera, 3 Microphones, F/T Sensor Franka Emika Panda Reaserach 3 Haptic teleoperation TAS, MPL, AD, TIL
TABLE I: Datasets comparison. We compare several commonly used datasets based on the number of demonstrations, the number of verbs they contain, the sensors used during data collection, the robotic platform, the data collection method, and the tasks that can be learned from the dataset. The tasks include TAS (Temporal Action Segmentation), MPL (Motion Policy Learning), AD (Anomaly Detection), and TIL (Task Inversion Learning). We use "n/r" to denote information that is not reported by the works and "N/A" for cases where the category is not applicable.
@@ II Related Works
Existing datasets, such as the Furniture Benchmark [16], have made progress in addressing long-horizon tasks like furniture assembly. However, they lack the high-quality force-torque data required for the tight tolerances demanded in these applications. Similarly, while datasets like RH20T include manipulation data with audio and force-torque sensor information, they do not address the long-horizon task sequences needed for complex assemblies [18]. Vision-based tactile datasets, such as SPARSH, provide valuable insights into contact-rich manipulation, but they too fall short in capturing the extended temporal dynamics and multi-modal sensing required for precise assembly and disassembly tasks [26]. To address these limitations, REASSEMBLE introduces a novel dataset incorporating high-resolution force-torque sensing specifically tailored for tight-tolerance, high-precision, and long-horizon tasks.
@@
The increasing prevalence of automation in robotic manipulation tasks highlights the necessity of effective skill assessment, task monitoring, and summarization to enhance system performance and reliability. Temporal action segmentation (TAS) plays a critical role in achieving these objectives by identifying and classifying distinct actions within continuous temporal sequences. Numerous datasets have been developed to support temporal action segmentation [20, 27, 28]. For example, the 50Salads dataset is a widely used benchmark where humans perform salad preparation tasks [20]. This dataset provides temporally labelled actions for long-duration videos, facilitating the training and evaluation of models for action segmentation. These models can learn to segment complex temporal sequences into meaningful sub-tasks. However, such datasets primarily focus on human activity and often lack relevance to robotic manipulation tasks. Robotics-specific datasets, such as RT-1, Bridge V2, and RoboSet, offer data tailored to specific robotic tasks [29, 15]. While these datasets are valuable for advancing task-specific learning, they are not designed for comprehensive temporal action segmentation across diverse manipulation contexts. Moreover, conventional action segmentation datasets like 50Salads predominantly rely on visual and temporal information, often overlooking critical multimodal data, such as force-torque measurements, which are essential for understanding robotic actions. Therefore, REASSEMBLE also addresses robotic action segmentation and incorporates multimodal data, including visual, force-torque, and temporal streams.
@@
Advancements in large vision-language models (LVLMs), such as PaLM-E, LLaVa, have shown impressive capabilities in reasoning about task conditions by leveraging visual information during robotic task execution [8, 30]. These models provide robust insights into system states, enabling robots to monitor and adapt to dynamic conditions. However, their significant computational requirements make them impractical for real-time applications, where lightweight and efficient models are essential for rapid decision-making. To address these challenges, ConditionNet was introduced as a focused resource for condition monitoring in robotic systems [23]. While ConditionNet dataset ((Im)PerfectPour) is not aimed at contact-rich manipulation and long-horizon task. Building on these limitations, REASSEMBLE is designed to address the gaps in existing resources. It includes conditioned labeled multi-modal data such as visual, force-torque, and temporal streams, enabling more comprehensive condition monitoring. Crucially, it incorporates failure data to train models that can effectively learn to detect, understand, and respond to failures in real time.
@@ III-A REASSEMBLE Tasks
The REASSEMBLE dataset encompasses both assembly and disassembly processes, with the associated actions defined as follows: for assembly, the actions are picking and inserting, whereas for disassembly, the actions are removing and placing. These four actions represent the minimal set required to comprehensively describe the assembly and disassembly tasks. The manual labelling of demonstrations by human operators necessitated limiting the action set to these four categories, thereby reducing cognitive load during annotation. The assembly board contains a total of 20 objects; however, due to constraints in the robot and teleoperation setup, the three nuts were excluded from the dataset (M12, M8, and M4). As a result, the dataset comprises 17 objects. A detailed breakdown of the objects included in the dataset is provided in the Appendix in Figure 13.
@@ III-B Sensors
In addition to the aforementioned sensors, the dataset includes an event camera. The sensor is mounted externally on a tripod in front of the robot. Due to its limited resolution and field-of-view, it is positioned at a distance that allows it to capture the majority of the workspace of the robot while still being able to perceive the smaller objects, such as small gears or pegs.
@@
Event cameras are emerging as critical sensors in robotics, providing a fundamentally different data acquisition paradigm compared to traditional RGB cameras [32]. Unlike frame-based cameras that capture static images at fixed intervals, event cameras asynchronously record changes in pixel intensity, resulting in high temporal resolution and low latency. Event cameras can provide valuable information for robotic manipulation by capturing motion events with exceptional precision. For instance, during manipulation tasks, event cameras can highlight fine-grained movements, such as the subtle shifts of objects in the environment. An example of this is shown in Figure 1, where the robot nudges a stuck peg. Once dislodged, the peg’s movement is clearly captured in the event data. Event camera information has also been successfully used in tasks such as slip detection [33], and has shown improved performance in human action recognition compared to using only RGB images [34]. Motivated by these findings, we incorporated event camera data into REASSEMBLE’s multi-modal framework.
(a) Before nudge.
(b) After nudge.
Figure 1: Visualization of event camera data. In this example, a peg becomes stuck after insertion, and the robot applies a nudge to properly insert it. (a) shows a snapshot of the event stream before the nudge, and (b) after. The motion of the peg is clearly visible in the event camera stream.
@@ III-C Teleoperation Setup
where $\mathbf{x}_{r}^{\text{init}}\in\mathbb{R}^{3}$ is the initial position of the robot’s end effector. The relative orientation change between the master’s orientation $\mathbf{q}_{m}=[w_{m},\mathbf{v}_{m}^{T}]\in\mathbb{R}^{4}$ and its initial orientation $\mathbf{q}_{m}^{\text{init}}\in\mathbb{R}^{4}$ is calculated as
@@
Given this quaternion difference $\Delta\mathbf{q}_{m}$ and the robot’s initial orientation $\mathbf{q}_{s}^{\text{init}}$, the target orientation of the robot end effector is computed as:
@@
The control input $\tau_{\rm{imp}}\in\mathbb{R}^{7}$ for the robot’s Cartesian impedance controller is expressed as
$\tau_{\rm{imp}}=\mathbf{J}^{\top}(\mathbf{K}_{x}\mathbf{e}_{\xi}+\mathbf{D}\mathbf{e}_{\dot{\xi}}),$ (5)
where $\mathbf{e}_{\xi}\in\mathbb{R}^{6}$ is the pose error vector derived from $\boldsymbol{\xi}_{\mathrm{r,cmd}}-\boldsymbol{\xi}_{\mathrm{r}}$, and $\mathbf{e}_{\dot{\xi}}=\dot{\boldsymbol{\xi}}_{\mathrm{r,cmd}}-\dot{\boldsymbol{\xi}}_{\mathrm{r}}\in\mathbb{R}^{6}$ is the twist error. Note that while $\boldsymbol{\xi}_{\mathrm{r}}\in\mathbb{R}^{7}$ and $\boldsymbol{\xi}_{\mathrm{r,cmd}}\in\mathbb{R}^{7}$ are the current and commanded Cartesian poses (with position $\mathbf{x}_{\mathrm{r}}\in\mathbb{R}^{3}$ and quaternion orientation $\mathbf{q}_{\mathrm{r}}\in\mathbb{R}^{4}$), the pose error $\mathbf{e}_{\xi}$ is computed as a 6D vector by appropriate transformation of the orientation difference. The matrices $\mathbf{K}_{x}\in\mathbb{R}^{6\times 6}$ and $\mathbf{D}\in\mathbb{R}^{6\times 6}$ represent the stiffness and damping coefficients, respectively.
@@
$V(\boldsymbol{\theta})=\alpha\det\left(\mathbf{J}(\boldsymbol{\theta})\mathbf{J}^{\top}(\boldsymbol{\theta})\right)+\beta\mathrm{J_{\rm L}}(\boldsymbol{\theta})+\gamma\dot{\boldsymbol{\theta}},$ (9)
where $\alpha,\beta,\gamma\in\mathbb{R}$ are weighting coefficients. The determinant term $\det\left(\mathbf{J}(\boldsymbol{\theta})\mathbf{J}^{\top}(\boldsymbol{\theta})\right)$ promotes the avoidance of singularities by maximizing the manipulability measure. The joint limit avoidance term $\mathrm{J_{\rm L}}(\boldsymbol{\theta})$ is expressed as
$\mathrm{J_{\rm L}}(\boldsymbol{\theta})=\tanh\left(a(\boldsymbol{\theta}-\mathbf{c})+b\right)+\tanh\left(-a(\boldsymbol{\theta}-\mathbf{c})+b\right)+2,$ (10)